OpenStudy (anonymous):

Solve: \[(2xy)dx + (y^2-3x^2)dy=0\]

OpenStudy (sirm3d):

\[\frac{ \partial }{ \partial y }(2xy)=2x,\frac{ \partial }{ \partial x }(y^2-3x^2)=-6x\]\[\frac{ -(\frac{ \partial M }{ \partial y } - \frac{ \partial N }{ \partial x })}{ 2xy }=-\frac{ 4 }{ y }\]the integrating factor that will make the DE exact is \[\large e^{\int\limits_{}^{}\frac{ -4 }{ y }dy}=\frac{ 1 }{ y^4 }\]

OpenStudy (sirm3d):

if you have not reached this part of the course, an alternate solution is \[\large x=vy\] and \[\large dx = vdy + ydv\] where the original DE is now separable in v and y.

OpenStudy (sirm3d):

have you solved this problem @RolyPoly ?

OpenStudy (anonymous):

I'm on the way solving it.. Sorry, please give me some more time...

OpenStudy (anonymous):

I'm not quite sure if I have learnt this way before. I have no memory of seeing this, but it appears in my book..

OpenStudy (sirm3d):

homogeneous function, doesn't it ring a bell?

OpenStudy (anonymous):

Nope :(

OpenStudy (sirm3d):

my comment on your attached file: \[\huge F=\frac{ x^2 }{ y^3 }+g_1(y)\] NOT \[\huge F_x\] and \[\huge F=-\frac{1}{y}+\frac{x^2}{y^3}+g_2(x)\] instead of \[\huge F_y\]

OpenStudy (anonymous):

:O That's how my teacher writes... \(F_x\) means integrating w.r.t. x, I think :S

OpenStudy (sirm3d):

about that alternate solution \[\large \deg(2xy)=2;\deg(y^2)=2;\deg(-3x^2)=2\] the degrees of the terms are identical. a substitution \[\large y = ux \text{ or }x=vy\] will transform the DE that is solvable by separation

OpenStudy (sirm3d):

taking \[x=vy \text{ and } dx=vdy+ydv\]\[\large 2(vy)y(vdy+ydv)+(y^2-v^2y^2)dy=0\]

OpenStudy (anonymous):

I suppose it is separable, but somehow, I can't separate it. I guess I have to check my calculations again first :/

OpenStudy (sirm3d):

collecting terms of dy and dv, \[\large (2v^2y^2+y^2-v^2y^2)dy+2vy^3dv=0\]or\[\large y^2(v^2+1)dy+2vy^3dv=0\]divide by \[\large y^3(v^2+1)\] and the DE becomes \[\large \frac{dy}{y}+\frac{2vdv}{v^2+1}=0\]

OpenStudy (anonymous):

How come... you can get it but I can't... :'(

OpenStudy (sirm3d):

expand after replacing x by vy and dx by vdy + ydv collect terms of dy and dv, simplify and factor.

OpenStudy (anonymous):

I tried w = y/x but it doesn't seem to be working for me :(

OpenStudy (anonymous):

Sorry, I got to go now.. I'd keep trying until I got the answer. Thanks for your help!!

OpenStudy (anonymous):

@sirm3d You're right about the F and Fx thing! Sorry for my earlier mistakes!!! When I tried w=y/x, I got something quite ugly :S